[Hallc_running] [New Logentry] RC daily update

sjjooste at jlab.org sjjooste at jlab.org
Fri Oct 12 14:20:03 EDT 2018

Logentry Text:
Before we start the actual report, I want to bring it to your attention that we still NEED A TARGET OPERATOR. for THIS SUNDAY’s (October 14) DAY SHIFT. We would greatly appreciate it if someone could help us out by signing up for this shift. Thank you in advance!

When running, beam quality was maybe a bit better today than yesterday, but we had 7 (seven!) major downtime events. You can find more details here: https://logbooks.jlab.org/entry/3608241
The big (3+h ongoing) downtime today is to address the high BLM trip rate. Jay Benesh stated this morning that the goal is to reduce the overal trip rate for the weekend by a factor of 2.

The end of the day shift and most of swing shift was able to take date relatively smoothly at 40uA (scheduled luminosity for this experiment is 70uA). 

Near the end of swing shift the SHSM HB QD alarm went off. We reset the QD https://logbooks.jlab.org/entry/3607947 which solved the issue and ramped back up to nominal current with the usual procedure https://logbooks.jlab.org/entry/3607952 
Note that might be storm related (happened at the same time the accelerator was down due to the storm). 

At the start of owl we were still without beam, so we rotated the SHMS to 6.495 degrees for the next kinematic point https://logbooks.jlab.org/entry/3607965
We encountered no issues. I would like to thank Holly for being on-call in case we tripped the rotation alarm (luckily we did not).

When beam was restored around 1am we noticed during the ramp-up that we were already hitting 700kHz singles rate in the SHMS at 35uA, yielding a 80% tracking efficiency. We briefly requested 30uA to see the impact on tracking efficiency. We found that at 30uA (15% less luminosity) the tracking efficiency only increased to 83%, so we went back to 35uA.

At 04:18, in the middle of owl, the BSY3C upstream dipole dropped, triggering a PSS alarm. All the run safe boxes dropped at the same time, and a mismatch between A-chain and B-chain caused the hall to be placed in restricted access. This led to 2h35min of lost beam time because a full RadCon survey was required before we could go back to beam permit https://logbooks.jlab.org/entry/3608094.

During the ongoing beam studies to address the BLM trip rate, MCC is sending currents up to 60uA to Hall C, which caused the SHMS DC1 chambers to trip (we are at a very small angle). We decided to leave the DC1 HV off until the beam studies are complete. https://logbooks.jlab.org/entry/3608212

We finished taking data in the Q2=3. W=3.14 x=0.25 theta_Shms=9.42 deg. We collected 46% of the statistics goal, short of the 50-60% we were hoping for, due to all the shutdowns after 11pm. We moved to the next kinematic point at Q2=3. W=3.14 x=0.25 theta_Shms=6.495 deg and proceeded to take a very meager 2% of the statistics goal until now. Note that this angle is slightly lower than what is in the printed run-plan for a more optimal kinematic coverage.

There will be an RC meeting today at 4:30 pm on the second floor of the counting house.

To join the Meeting:

To join via Room System:
Video Conferencing System: bjn.vc -or-
Meeting ID : 861439920

To join via phone :
1) Dial:
+1.408.740.7256 (US (San Jose))
+1.888.240.2560 (US Toll Free))
+1.408.317.9253 (US (Primary, San Jose))
(see all numbers - http://bluejeans.com/numbers)

2) Enter Conference ID : 861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3608253
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181012/eac57cc5/attachment.html>

More information about the Hallc_running mailing list